# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

### v1

@@ Abstract
Expensive multi-objective optimization is a prevalent and crucial concern in many real-world scenarios, where sample-efficiency is vital due to the limited evaluations to recover the true Pareto front for decision making. Existing works either involve rebuilding Gaussian process surrogates from scratch for each objective in each new problem encountered, or rely on extensive past domain experiments for pre-training deep learning models, making them hard to generalize and impractical to cope with various emerging applications in the real world. To address this issue, we propose a new paradigm named FoMEMO (Foundation Models for Expensive Multi-objective Optimization), which enables the establishment of a foundation model conditioned on any domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregation posteriors. Rather than accessing extensive domain experiments in the real world, we demonstrate that pre-training the foundation model with a diverse set of hundreds of millions of synthetic data can lead to superior adaptability to unknown problems, without necessitating any subsequent model training or updates in the optimization process. We evaluate our method across a variety of synthetic benchmarks and real-word applications, and demonstrate its superior generality and competitive performance compared to existing methods.
@@ 1 Introduction
Optimizing multiple competitive objectives simultaneously is ubiquitous in many real-world applications, such as neural architecture search (Ying et al. 2019), engineering design (Tanabe and Ishibuchi 2020) and scientific discovery (Dara et al. 2022). The primary difficulty in these scenarios stems from the black-box nature of real-world evaluations, which are often costly and time-consuming. When decision-makers have access to only a limited number of evaluations, it becomes challenging to accurately recover the unknown true Pareto fronts.
@@
To achieve sample-efficiency, traditional approaches, such as multi-objective Bayesian optimization, typically build Gaussian process (GP) models (Williams and Rasmussen 1995) as the surrogates to approximate the objective functions, and then optimize well-designed acquisition functions to generate candidate solutions continuously to update the surrogates in an online manner (Daulton, Balandat, and Bakshy 2020, 2021). However, these methods typically need to rebuild surrogates from scratch for each objective when encountering a new problem, and the efficiency is greatly limited by the cost of GP training and inference (Williams and Rasmussen 2006). To enhance the generalization and scalability in downstream optimization, some research focuses on pre-training deep learning models in offline scenarios (Xue et al. 2024; Yuan et al. 2024). These approaches typically assume accessing to a large amount of real-world experimental data is available for training. However, acquiring large datasets in practical applications, especially in unknown or emerging fields, is often challenging or even impossible. Moreover, these methods also require training dedicated models for specific problems, and hardly be generalized to unseen scenarios.
@@
In real-world contexts, a wide range of multi-objective optimization problems arise daily across various fields, each exhibiting unique characteristics. Given the limitations of existing methodologies, a compelling research question arises: Can we devise a universal method that can efficiently address a wide range of emerging real-world multi-objective optimization problems?
@@
In this paper, we attempt to answer this question by exploring a new paradigm named FoMEMO. It enables us to establish a foundation model conditioned on arbitrary domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregation posteriors. What distinguishes our approach from existing methods lies in two aspects. Firstly, unlike traditional online methods that require rebuilding a surrogate from scratch for each objective in newly encountered problems and necessitate frequent, costly model updates during optimization, our approach involves pre-training a foundation model only once, which can be directly adapted to arbitrary scenarios throughout the entire optimization process in an in-context manner. This means that by simply providing domain trajectories and user preferences as contextual information, the pre-trained model can achieve few-shot generalization in unknown scenarios to initiate fast optimization without the need for any subsequent model training or updates. Secondly, rather than accessing expensive real-world data from extensive past experiments, we continuously sample a large amount of diverse synthetic data from a computationally efficient sampling mechanism during pre-training. By simulating a large set of potential scenarios that may exist in the real world, the foundation model emerges with the ability to learn to predict the aggregated posteriors conditioned on any contexts.
@@
  • We pre-train a foundation model using hundreds of millions of sampled synthetic data, enabling it to learn to predict aggregated posteriors conditioned on arbitrary domain trajectories and user preferences. This approach eliminates the need for access to expensive experimental data from the real world, allowing for efficient and adaptable optimization across diverse scenarios.
@@
  • We evaluate the proposed method on both synthetic problems and real-world applications. The experimental results show that the proposed method demonstrates superior adaptability and generality across a wide range of unseen problems. In comparison to existing methods, our approach exhibits competitive optimization performance and high efficiency in the majority of cases.
@@ 2.1 Multi-objective Bayesian Optimization
Multi-objective Bayesian optimization (MOBO) aims to identify a finite set of Pareto optimal solutions within a constrained evaluation budget in an online manner. A popular class of MOBO algorithms is based on decomposition, wherein the original MOP is converted into several single-objective sub-problems that are easier to solve. For example, ParEGO (Knowles 2006) and TS-TCH (Paria, Kandasamy, and Póczos 2020) randomly select a preference in each iteration and optimize the corresponding sub-problem using single-objective BO. MOEA/D-EGO (Zhang et al. 2009) pre-defines several sub-problems with uniformly distributed preferences and then employs the population-based MOEA/D algorithm (Zhang and Li 2007) to jointly optimize the utilities of all sub-problems. Another representative type of MOBO algorithm is based on indicators. For instance, (Emmerich, Giannakoglou, and Naujoks 2006) developed expected hypervolume improvement (EHVI), which extends from the widely used expected improvement (EI) acquisition function (Jones, Schonlau, and Welch 1998), providing an overall performance across all objectives with uncertainty. To extend EHVI to the parallel evaluation setting, qEHVI (Wada and Hino 2019; Daulton, Balandat, and Bakshy 2020) was developed and further adapted to noisy scenarios (Daulton, Balandat, and Bakshy 2021). In addition to the application in designing acquisition functions, multi-objective indicators have also been utilized in batch sample selection before expensive evaluations (Bradford, Schweidtmann, and Lapkin 2018; Konakovic Lukovic, Tian, and Matusik 2020; Lin et al. 2022). While existing methods have demonstrated their effectiveness, it is important to note that they require the modeling of each objective for every new problem from the ground up. Additionally, the costs associated with GP training and inference further constrain their scalability.
2.2 Offline Multi-objective Optimization
@@
Beyond online optimization, there also exit scenarios where only pre-collected offline data is accessible and iterative online evaluations are not permitted. Although our work mainly focuses on online methods, we still present a briefly introduction here to highlight the superiority of the synthetic training paradigm in our method. In the offline context, a common approach involves training a deep neural network (DNN) on a static dataset to serve as an oracle (Trabucco et al. 2021; Yu et al. 2021; Chen et al. 2023). This is followed by applying search methods, such as gradient descent or evolutionary algorithms, to generate a set of candidate solutions that are potentially superior to those present in the dataset. (Xue et al. 2024) introduced the first benchmark for offline multi-objective optimization and proposed DNN-based and GP-based methods. (Yuan et al. 2024) developed ParetoFlow, specifically designed to guide flow sampling to approximate the Pareto front. However, existing offline methods typically require a large amount of training data from expensive experiments, which is often impractical for various emerging applications in the real-world. Moreover, these methods often necessitate the training of dedicated models for specific problems, which limits their ability to generalize to unseen scenarios.
@@ 2.3 Universal Regression Based on Transformers
Recently, Transformers (Vaswani et al. 2017) have become the preferred architecture for deep learning and foundation models. As model sizes and training data scaling, the models exhibit remarkable in-context learning (ICL) capabilities (Brown et al. 2020): they can make accurate predictions quickly in new tasks when prompted with only a few training examples, without any parameter update. Several research has demonstrated that Transformers can learn in-context simple logistic regression algorithms (Garg et al. 2022; Akyürek et al. 2022). Beyond predicting labels, Transformer Neural Processes (Nguyen and Grover 2022) enhance the prediction with uncertainty as an alternative framework for uncertainty-aware meta learning. Prior-data Fitted Networks (PFNs) (Müller et al. 2021) further show that Transformers can even learn to perform Bayesian inference with ICL. The developed universal regression models not only provide new insights into various model-based prediction topics (Hollmann et al. 2022, 2025; Dooley et al. 2023; Song et al. 2024a), but also promote the regress-then-optimize framework towards a generalized optimization paradigm in both numerical (Müller et al. 2023; Nguyen, Agrawal, and Grover 2023; Rakotoarison et al. 2024) and even linguistic space (Nguyen et al. 2024; Tan et al. 2025). Building on this potential, we aim to develop a general optimization method for multi-objective scenarios, aligning more closely with real-world challenges that involve multi-criteria outputs.
@@ 3.2 Scalarization for Multiple Objectives
where $y_{i}(\boldsymbol{x})$ is the $i$-th observation value of the corresponding true function $f_{i}(\boldsymbol{x})$ at the input $\boldsymbol{x}$, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})$ is the ideal point (lower bound) in the objective space. A promising property behind Tchebycheff scalarization function is that we can identify each Pareto optimal solution by optimizing the scalarization function with a specific but unknown preference (Choo and Atkins 1983):
$\boldsymbol{x}^{*}=\arg\min_{\boldsymbol{x}\in\mathcal{X}}s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})).$ (3)
In out work, we use the negative of the scalarization function as the aggregation target $g_{\boldsymbol{\lambda}}(\boldsymbol{x})=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ for the convenience of model training and inference. Thus, we wish to maximize $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given each preference vector $\boldsymbol{\lambda}$.
Another representative scalarization method is the hypervolume (HV) indicator (Zitzler and Thiele 2002), which is monotonic to the Pareto dominance relation. The hypervolume of a finite approximate Pareto front $\mathcal{P}$ is defined as:
@@ 3.3 Bayesian Optimization
Bayesian optimization (BO) (Shahriari et al. 2015; Frazier 2018; Garnett 2023) is a sample-efficient technique for solving expensive black-box optimization problems. BO utilizes probabilistic surrogate models, typically Gaussian processes (GPs) (Williams and Rasmussen 2006), to provide an prediction with calibrated uncertainty for the latent functions $\boldsymbol{f}$ based on the true observed data $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$. To suggest the candidate solutions for next evaluations, BO defines and optimizes an acquisition function $\alpha:\mathcal{X}\mapsto\mathbb{R}$, which quantifies the utility value of evaluating a new set of solutions based on the predictive distribution of the surrogate models, effectively balancing exploration and exploitation. BO continuously generates candidate solutions for evaluation and updates the surrogate models in an online manner until the total evaluation budget is exhausted.
@@ 4 Methodology
Figure 1: Overview of FoMEMO framework, including synthetic pre-training stage (left) and in-context optimization stage (right). During synthetic pre-training, the foundation model is fed with a large set of synthetic data as the context inputs, which include the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with its aggregation label $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the corresponding artificial preference $\boldsymbol{\lambda}$ sampled on the simplex. The model parameterized by $\boldsymbol{\theta}$ is trained once only to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ as the output conditioned on the contextual information. During the in-context optimization stage, users can readily address new problems by simply providing evaluated trajectories from arbitrary unseen real-world applications, along with any potential user preferences (if applicable) as prompt inputs for the foundation model. Using the predicted aggregated posteriors as the basis, various acquisition functions can be naturally derived to enable fast in-context optimization. These acquisition functions can be efficiently optimized to suggest the next candidate within the utility of interest, free from any additional model training or updates throughout the entire optimization process.
@@ 4.1 Framework
Instead of building a separate surrogate model from scratch for each objective function in every new problem encountered, the core concept of our method is to develop a universal foundation model. This model is pre-trained once only using a vast corpus of synthetic samples and preferences as context information, and outputs a large set of preference-wise aggregation posterior distributions for fast in-context optimization. We present the details of our framework in Figure 1, which includes synthetic pre-training stage on the left and in-context optimization stage on the right.
Specifically, in the synthetic pre-training stage, rather than accessing expensive real-world experiments, we implement a data sampling mechanism that allows us to heavily sample from it to generate a large set of synthetic data as the context inputs, including the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with the corresponding aggregation label $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the artificial preference $\boldsymbol{\lambda}$ sampled on the simplex, constructing a supervised learning problem for training the foundation model efficiently. We repeatedly synthesize training data as above and optimize the model’s parameters $\boldsymbol{\theta}$ to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ conditioned on the contextual inputs. We will present the details of data generation in Section 4.2.
Once the pre-training is completed, during the in-context optimization stage, users can tackle new problems by simply providing the evaluated trajectory from arbitrary unseen real-world applications, as well as the the user’s potential preferences (if any) as prompt inputs for the foundation model. Based on the aggregated posteriors predicted by the foundation model in a single neural network forward pass, various acquisition functions can be easily derived to initiate fast in-context optimization to generate the next candidate for expensive evaluation for the utility of interest, without the need for any additional model training or updates during the optimization process.

### v2

@@ Abstract
Expensive multi-objective optimization is a prevalent and crucial concern in many real-world scenarios, where sample-efficiency is vital due to the limited evaluations to recover the true Pareto front for decision making. Existing works either involve rebuilding Gaussian process surrogates from scratch for each objective in each new problem encountered, or rely on extensive past domain experiments for pre-training deep learning models, making them hard to generalize and impractical to cope with various emerging applications in the real world. To address this issue, we propose a new paradigm named FoMEMO (Foundation Models for Expensive Multi-objective Optimization), which enables the establishment of a foundation model conditioned on any domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregated posteriors. Rather than accessing extensive real-world domain experiments for training, we demonstrate that pre-training the foundation model with a diverse set of hundreds of millions of synthetic data can lead to superior generalization and optimization performance to unknown problems, without necessitating any subsequent model training or updates in the following optimization process.
Machine Learning, ICML
@@ 1 Introduction
Optimizing multiple competitive objectives simultaneously is ubiquitous in many real-world applications, such as neural architecture search (Ying et al., 2019), engineering design (Tanabe and Ishibuchi, 2020) and scientific discovery (Dara et al., 2022). The primary difficulty in these scenarios stems from the black-box nature of real-world evaluations, which are often costly and time-consuming. When decision-makers have access to only a limited number of evaluations, it becomes challenging to accurately recover the unknown true Pareto fronts.
@@
To achieve sample-efficiency, traditional multi-objective Bayesian optimization (MOBO) typically employs Gaussian Process (GP) surrogates (Williams and Rasmussen, 1995) to approximate objective functions, and then optimize acquisition functions to generate candidate solutions to update the surrogates iteratively in an online manner (Daulton et al., 2020, 2021). However, these methods typically need to rebuild surrogates from scratch for each objective when encountering a new problem, and the efficiency is greatly limited by the cost of GP training and inference (Williams and Rasmussen, 2006). To enhance the generalization and scalability in downstream optimization, some research focuses on pre-training deep learning models in offline scenarios (Xue et al., 2024; Yuan et al., 2024). These approaches typically assume accessing to a large amount of real-world experimental data is available for training. However, acquiring large datasets in practical applications, especially in unknown or emerging fields, is often challenging or even impossible. While recent work has explored pre-training with synthetic functions in the online setting (Hung et al., 2025), its architecture still depends heavily on the repetitive GP modeling. Furthermore, the reliance on computationally intensive hypervolume-based rewards (Zitzler and Thiele, 2002) imposes a significant bottleneck on training scalability, thereby restricting the diversity of the training data and hindering its generalization to unseen scenarios.
@@
In real-world contexts, a wide range of multi-objective optimization problems arises daily across various fields, each exhibiting unique characteristics and demand rapid solutions. Given the limitations of existing methodologies, a compelling research question arises:
Can we devise a universal method that can efficiently address a wide range of emerging real-world multi-objective optimization problems?
@@
In this paper, we attempt to answer this question by exploring a new paradigm named FoMEMO. It enables us to establish a foundation model conditioned on an arbitrary domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregated posteriors. What distinguishes our approach from existing methods lies in two aspects. Firstly, unlike traditional MOBO methods that require rebuilding a surrogate from scratch for each objective in newly encountered problems and necessitate frequent, costly model updates during optimization, our approach involves pre-training a foundation model only once, which can be directly adapted to arbitrary scenarios throughout the entire optimization process in an in-context manner. This means that by simply providing domain trajectories and user preferences as contextual information, the pre-trained model can achieve few-shot generalization in unknown scenarios to initiate fast optimization without any subsequent model training or updates. Secondly, rather than accessing expensive real-world data from extensive past experiments for training, we continuously sample a large amount of diverse synthetic data from a computationally efficient sampling mechanism during pre-training. By simulating a large set of potential scenarios that may exist in the real world, the foundation model emerges with the ability to learn to predict the aggregated posteriors conditioned on any contexts.
@@
  • We pre-train a foundation model using hundreds of millions of sampled synthetic data, enabling it to learn to predict aggregated posteriors conditioned on arbitrary domain trajectories and user preferences. This approach eliminates the need for access to extensive expensive experimental data from the real world during training, allowing for efficient and adaptable optimization across diverse scenarios.
@@
  • We evaluate the proposed method on a diverse set of synthetic problems and real-world applications. Experimental results demonstrate that the method exhibits remarkable adaptability and generalization capabilities across a wide range of unseen problems. In comparison to existing methods, our approach consistently achieves superior optimization performance and high efficiency in the majority of cases.
@@ 2.1 Multi-objective Bayesian Optimization
Multi-objective Bayesian optimization (MOBO) typically employs GP models and acquisition functions to identify Pareto optimal solutions within a constrained evaluation budget. A popular category of MOBO is based on decomposition, wherein the original MOP is converted into several single-objective sub-problems that are easier to solve (e.g., ParEGO (Knowles, 2006), TS-TCH (Paria et al., 2020) and MOEA/D-EGO (Zhang et al., 2009)). Besides decomposition, indicator-based approaches provide an integrated performance measure across all objectives under uncertainty. Prominent examples include EHVI (Emmerich et al., 2006), which generalizes the Expected Improvement (EI) function (Jones et al., 1998) to multi-objective settings, along with its modern variants: the parallelizable qEHVI (Wada and Hino, 2019; Daulton et al., 2020) and the noise-robust qNEHVI (Daulton et al., 2021). Another representative class stems from information theory, focusing on selecting points that maximize information gain regarding the Pareto set or front (e.g., PESMO (Hernández-Lobato et al., 2016), MESMO (Belakaria et al., 2019) and JES (Tu et al., 2022; Hvarfner et al., 2022). In addition to model management and acquisition functions design, various batch sample selection strategies have also been developed in parallel setting (e.g., TSEMO (Bradford et al., 2018), USEMO-EI (Belakaria et al., 2020) and DGEMO (Konakovic Lukovic et al., 2020)). While effective, these traditional MOBO methods typically require training individual models for each objective of every new problem from scratch. Furthermore, the substantial computational overhead associated with GP training and inference significantly hinders their scalability.
2.2 Learning-based Expensive Optimization
@@
To enhance the generalization and scalability in downstream optimization, several approaches in the offline context involves training a deep neural network (DNN) on a static dataset to serve as an oracle (Trabucco et al., 2021; Yu et al., 2021; Chen et al., 2023). This is followed by applying search methods, such as gradient descent or evolutionary algorithms, to generate a set of candidate solutions that are potentially superior to those present in the dataset (Xue et al., 2024; Yuan et al., 2024). However, these approaches typically require a large amount of training data from expensive experiments, which is often impractical for various emerging applications in the real-world. Moreover, they often necessitate the training of dedicated models for specific problems, which limits their ability to generalize to unseen scenarios. To circumvent the need for extensive real-world training data, inspired by meta-learning frameworks in single-objective BO (Volpp et al., 2019; Hsieh et al., 2021; Maraval et al., 2024), BOFormer (Hung et al., 2025) has recently explored pre-training policies via reinforcement learning on synthetic data in multi-objective context. However, to predict acquisition function values, its architecture remains heavily dependent on iterative GP modeling to generate posterior distributions. Furthermore, the computationally intensive nature of hypervolume-oriented rewards imposes a significant bottleneck on training scalability, thereby restricting the diversity of the pre-training corpus and hindering its generalization to unseen complex problems.
@@ 2.3 Universal Regression Based on Transformers
Recently, Transformers (Vaswani et al., 2017) have become the preferred architecture for deep learning and foundation models. As model sizes and training data scaling, the models exhibit remarkable in-context learning (ICL) capabilities (Brown et al., 2020): they can make accurate predictions quickly in new tasks when prompted with only a few training examples, without any parameter update. Several research has demonstrated that Transformers can learn in-context simple logistic regression algorithms (Garg et al., 2022; Akyürek et al., 2022). Beyond predicting labels, Transformer Neural Processes (Nguyen and Grover, 2022) enhance the prediction with uncertainty as an alternative framework for uncertainty-aware meta learning. Prior-data Fitted Networks (PFNs) (Müller et al., 2021) further show that Transformers can even learn to perform Bayesian inference with ICL. The developed universal regression models not only provide new insights into various model-based prediction topics (Hollmann et al., 2022, 2025; Dooley et al., 2023; Song et al., 2024a), but also promote the regress-then-optimize framework towards a generalized optimization paradigm in both numerical (Müller et al., 2023; Nguyen et al., 2023; Rakotoarison et al., 2024) and even linguistic space (Nguyen et al., 2024; Tan et al., 2025). Building on this potential, we aim to develop a general optimization method for multi-objective scenarios, aligning more closely with real-world challenges that involve multi-criteria outputs.
@@ 3.2 Scalarization for Multiple Objectives
where $y_{i}(\boldsymbol{x})$ is the $i$-th observation value of the corresponding true function $f_{i}(\boldsymbol{x})$ at the input $\boldsymbol{x}$, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})$ is the ideal point (lower bound) in the objective space. A promising property behind Tchebycheff scalarization function is that we can identify each Pareto optimal solution by optimizing the scalarization function with a specific but unknown preference (Choo and Atkins, 1983):
$\boldsymbol{x}^{*}=\arg\min_{\boldsymbol{x}\in\mathcal{X}}s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x}).$ (3)
In our work, we use the negative of the scalarization function as the aggregation target $g_{\boldsymbol{\lambda}}(\boldsymbol{x})=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ for the convenience of model training and inference. Thus, we wish to maximize $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given each preference vector $\boldsymbol{\lambda}$.
Another representative scalarization method is the hypervolume (HV) indicator (Zitzler and Thiele, 2002), which is monotonic to the Pareto dominance relation. The hypervolume of a finite approximate Pareto front $\mathcal{P}$ is defined as:
@@ 3.3 Bayesian Optimization
Bayesian optimization (BO) (Shahriari et al., 2015; Frazier, 2018; Garnett, 2023) is a sample-efficient technique for solving expensive black-box optimization problems. BO utilizes probabilistic surrogate models, typically Gaussian processes (GPs) (Williams and Rasmussen, 2006), to provide an prediction with calibrated uncertainty for the latent functions $\boldsymbol{f}$ based on the true observed data $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$. To suggest the candidate solutions for next evaluations, BO defines and optimizes an acquisition function $\alpha:\mathcal{X}\mapsto\mathbb{R}$, which quantifies the utility value of evaluating a new set of solutions based on the predictive distribution of the surrogate models, effectively balancing exploration and exploitation.
@@ 4 Methodology
Figure 1: Overview of FoMEMO framework, including synthetic pre-training stage (left) and in-context optimization stage (right). During synthetic pre-training, the foundation model is fed with a large set of synthetic data as the context inputs, which include the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with its aggregation target $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the corresponding artificial preference $\boldsymbol{\lambda}$ sampled on the simplex. The model parameterized by $\boldsymbol{\theta}$ is trained once only to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ as the output conditioned on the contextual information. During the in-context optimization stage, users can readily address new problems by simply providing evaluated trajectories from arbitrary unseen real-world applications, along with any potential user preferences (if applicable) as prompt inputs for the foundation model. Using the predicted aggregated posteriors as the basis, various acquisition functions can be naturally derived to enable fast in-context optimization. These acquisition functions can be efficiently optimized to suggest the next candidate within the utility of interest, free from any additional model training or updates throughout the entire optimization process.
@@ 4.1 Framework
Instead of building a separate surrogate model from scratch for each objective function in every new problem encountered, the core concept of our method is to develop a universal foundation model. Our foundation model is pre-trained once only using a vast corpus of synthetic samples and preferences as context information, and outputs a large set of preference-wise aggregated posterior distributions for fast in-context optimization. We present the details of our framework in Figure 1, which includes a synthetic pre-training stage on the left and an in-context optimization stage on the right.
Specifically, in the synthetic pre-training stage, rather than accessing expensive real-world experiments, we implement a data sampling mechanism that allows us to heavily sample from it to generate a large set of synthetic data as the context inputs, including the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with the corresponding aggregation target $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the artificial preference $\boldsymbol{\lambda}$ sampled on the simplex, constructing a supervised learning problem for training the foundation model efficiently. We repeatedly synthesize training data as above and optimize the model’s parameters $\boldsymbol{\theta}$ to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ conditioned on the contextual inputs. We will present the details of data generation in Section 4.2.
Once the pre-training is completed, during the in-context optimization stage, users can tackle new problems by simply providing the evaluated trajectory from arbitrary unseen real-world applications, as well as the user’s potential preferences (if any) as prompt inputs for the foundation model. Based on the aggregated posteriors predicted by the foundation model in a single neural network forward pass, various acquisition functions can be easily derived to initiate fast in-context optimization to generate the next candidate for expensive evaluation for the utility of interest, without the need for any additional model training or updates during the optimization process.